OpenStudy (anonymous):

Would someone please help me with this? I just don't understand how to do it. Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

"the sum of the lateral area and area of one base was about 3000 square centimeters" this means that if we add up the surface area of the sides and base it = 3000 so, surface area of the sides = 2*pi*r*h and surface area of base = pi*r^2 this means that: 3000 = 2*pi*r*h + pi*r^2 we know that h= 50 so, 3000 = 100*pi*r + pi*r taking pi to be 3.14: 3000 = 314*r + 3.14r 3000 = 317.14*r r = 9.46 cm (estimation)

OpenStudy (anonymous):

oops wait i forgot a ^2 in there

OpenStudy (anonymous):

3000 = 100*pi*r + pi*r^2 taking pi to be 3.14: 3000 = 314*r + 3.14r^2 setting = 0 3.14r^2 + 314r -3000 = 0 we must use the quadratic formula: \[\frac{ -314 \pm \sqrt{314^{2}-4(3.14)(-3000)} }{ 2(3.14) }\]

OpenStudy (anonymous):

Umm What's a setting?

OpenStudy (anonymous):

which gives: r = 8.78 cm and r = -108.78 cm but a negative radius doesn't make sense so: r = 8.78 cm

OpenStudy (anonymous):

*You're setting the equation equal to 0 in that step

OpenStudy (anonymous):

But why?

OpenStudy (anonymous):

@wolfgirl I'm not sure how to do it either. Want to work on it together?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Okay. What class are you in? I mean how much do you know about Geometry?

OpenStudy (anonymous):

@wolfgirl are you pressed for time, or can you take your time?

OpenStudy (anonymous):

You there??

OpenStudy (anonymous):

I'm so sorry! I was in the middle of something. I'm here now.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I'm in this Geometry Class in Ashworth High School but it isn't my strongest subject.

OpenStudy (anonymous):

What year? What do you know about geometry? A little. A lot. Sort of okay.

OpenStudy (anonymous):

I'm in my senior year. I've basically gone through the entire class and I'm working on my finals. I know the most of the formulas and stuff.

OpenStudy (anonymous):

You know more than I do. What was it about Piglet's explanation that you didn't understand?

OpenStudy (anonymous):

Haha Well I don't understand what a setting is or what it's used for...

OpenStudy (anonymous):

Okay. Let me read what he wrote.

OpenStudy (anonymous):

Let's forget about Piglet. What's a lateral?

OpenStudy (anonymous):

Umm I believe that a lateral is the sum of all the faces

OpenStudy (anonymous):

I think you're going to need to be certain about that, aren't you? How much time do you have?

OpenStudy (anonymous):

I have plenty of time

OpenStudy (anonymous):

Okay. Maybe we should be working from the same textbook. I'll give you the link to some free ones.

OpenStudy (anonymous):

Okay, thank you

OpenStudy (anonymous):

http://opencoursescholar.wordpress.com/resources/openstax-textbooks/ Go down to geometry and click on that. Then when you're on the page, click on flexbooks.

OpenStudy (anonymous):

Okay. There

OpenStudy (anonymous):

There are two books that I recommend. One is Geometry Concepts and the other is Geometry 2

OpenStudy (anonymous):

I mean once you're on the Flexbooks page.

OpenStudy (anonymous):

Okay, I'm on Geometry Concepts

OpenStudy (anonymous):

Download it as a PDF

OpenStudy (anonymous):

That way you'll have it whenever you need it. There are link in the books that you can click on that will take you to videos, and widgets, and so forth.

OpenStudy (anonymous):

Don't worry if you're having any trouble. That always seems to happen to me. Maybe I can help if you are having any.

OpenStudy (anonymous):

I'm downloading it

OpenStudy (anonymous):

Okay cool. Once you get that one, get the Geometry 2. I also found a video that you might like. Are you familiar with the Khan Academy?

OpenStudy (anonymous):

I am not familiar with them

OpenStudy (anonymous):

Okay. Well, you have time to relax and find out more about your available resources as time goes by. I found out that as you mentioned, they probably are talking about "lateral surface area". http://ncalculators.com/area-volume/cylinder-calculator.htm

OpenStudy (anonymous):

The link that I gave you earlier is for my blog. You'll find tons of excellent stuff there. So place the link some place that's handy for you. https://www.google.com/url?sa=t&rct=j&q=&esrc=s&source=web&cd=1&cad=rja&ved=0CC4QFjAA&url=http%3A%2F%2Fopencoursescholar.wordpress.com%2F&ei=-zRxUosM08-wBITdgcAN&usg=AFQjCNH-TGaxkrFCTIdwjft_MAhpTj_qKQ&sig2=agqJZEGG5ZDzRw8TJqqzyw&bvm=bv.55617003,d.cWc

OpenStudy (anonymous):

http://opencoursescholar.wordpress.com/ Above is the short version of the same link

OpenStudy (anonymous):

Thank you Well the problem was basically the fact that the number 3000 was the lateral area plus the area of one of the bases and I needed to find the radius. I was supposed to use the quadratic formula to find it. I'm not sure what that is though.

OpenStudy (anonymous):

You'll be using the quadratic formula A LOT!!!!! Are you planning on going to college?

OpenStudy (anonymous):

uhh yeah, I guess so

OpenStudy (anonymous):

Well, even if you do independent studies, you'll need this stuff. Some folks don't have the funds for college, but these days you can still study it.

OpenStudy (anonymous):

Okay, thanks. If I go to college, my parents are paying for it because I don't wanna go...they're making me.

OpenStudy (anonymous):

You'll probably be glad you went. Beats working at McDonalds. :)

OpenStudy (anonymous):

Here's a link to the quadratic formula http://www.ck12.org/algebra/Quadratic-Formula/

OpenStudy (anonymous):

Haha, and thanks

OpenStudy (anonymous):

You're very welcome.

OpenStudy (anonymous):

So, where are you at with this problem? That one link really has good stuff. The one with the applet.

OpenStudy (anonymous):

I think I've figured it out, it's more simple than I thought.

OpenStudy (anonymous):

Okay. Anything else you need?

OpenStudy (anonymous):

Nope, that'll be it :) Thank you so much!

OpenStudy (anonymous):

You're welcome. Good luck with the test. :)

OpenStudy (anonymous):

Thanks, I'll need it.